# Dynamic Multi-Hop HARQ (DMH-HARQ): Reliable and Open Latency-Constrained Wireless Transport Network

### v3

@@ VI-C Results and Analyses
We observe from the results that the DMH-HARQ approach consistently outperforms all baselines, even when under an unfair disadvantage in soft combining. It is promised to hold even more performance gain over the static HARQ solutions when applied with a more advanced combining technology. This excellent performance is credited to the DP algorithm that guarantees to achieve the theoretical maximum of the Bellman sum (12). Especially, with more available blocklength per hop, since the overall possible numbers of HARQ attempts over the $I$-hop chain increases, this performance gain brought by DP is also increasing. To the contrary, when the channel is harsh or the radio resource is extremely limited, the constraint (III-C) is likely forcing to allow only one transmission slot per hop, so that the optimal DMH-HARQ schedule converges to the optimal one-shot-per-hop schedule, with just gives the performance upper bound of optimal static HARQ-IR in FBL [5].
Figure 3: Sensitivity tests regarding (3(a)) the time frame length, (3(b)) the number of hops, (3(c)) the decoding/feedback delay, and (3(d)) the pathloss, respectively.
Remark that the listening-based cooperative ARQ approach also allocates the radio resource to different hops in a technically dynamic manner. However, its mechanism design only allows the successor hops to inherit unused resource from predecessors, not the other way around. With a same fixed blocklength for every (re)transmission, it results in a longer effective sub-frame for the late hops that are closer to the final information sink than the for early hops that are closer to the source. However, as it can be derived from [37], as a stochastic DP, the optimal solution shall be just the opposite: to allocate more resource to the early hops than the late ones. In the end, the listening-based cooperative ARQ approach cannot hit the optimum of the DP problem (III-C), and has only a limited gain from the dynamic scheduling. This impact is even more critical in the FBL regime where ARQ/HARQ generally work poor, leading to a significant gap to the performance in the infinite blocklength (IBL) regime. As we see from the results, the listenting-based cooperative ARQ approach performs the worst among all baselines.
@@ VII Further Discussions
VII-A Computational Complexity
Concerning the recursive nature of Algorithm 1, its computational load can be a crucial factor in assessing its practical applicability. Since DMH-HARQ can be considered as an $I$-hop generalization of the $2$-hop CLARQ protocol, we can follow the approach we used in [33] to analyze its time complexity, which returns us a result of $\mathcal{O}\left(n_{\mathrm{MAX}}^{I}\right)$.
It must be noted that, though this time complexity is significantly higher than to optimize the schedule of conventional static HARQ approaches, i.e. $\mathcal{O}\left[\log\left(n_{\mathrm{MAX}}\right)\right]$, the optimizer is not necessarily executed in real time for every communication session. Indeed, the optimal schedules regarding different CSI can be obtained offline for every node with Algorithm 1, and cumulatively stored in a LUT. In the runtime, the node only needs to look up the table for its optimal decision, which has a minimal time complexity of $\mathcal{O}\left(1\right)$.
VII-B IR Gain
@@
We have proven that even when applied without any combining technique, DMH-HARQ is able to outperform the optimal static HARQ-IR. A natural idea, of course, is to further enhance the performance of DMH-HARQ with soft combining, especially with an ideal IR, where the PER formula (8) must be replaced by
$\varepsilon_{s}=Q\left(\sqrt{\frac{n_{s}+\frac{\tau_{s}}{T_{\mathrm{b}}}}{V_{I+1-i_{s}}}}\left(\mathcal{C}_{I+1-i_{s}}-\frac{d}{n_{s}+\frac{\tau_{s}}{T_{\mathrm{b}}}}\right)\ln 2\right).$ (14)
@@
Though IR is guaranteed to reduce the achievable E2E packet loss rate of DMH-HARQ, it costs more computational effort. The inclusion of $\tau_{s}$ into $\varepsilon_{s}$ rejects Lemma 1. To keep it an MDP so that the integer DP based framework of Algorithm 1 still applies, $\tau_{s}$ must play its role as part of the system state when solving the decision policy, i.e. $n_{s}$. The dictionaries $\mathcal{D}_{\mathrm{local}}$ and $\mathcal{D}_{\mathrm{succ}}$ in Algorithm 1 shall map $[t_{s},i_{s},\tau_{s}]$ instead of $[t_{s},i_{s}]$ to $[\xi_{\mathrm{opt}},n_{\mathrm{opt}}]$. This will significantly increase the effective size of the state space, raising the time complexity of solving $n_{s}$ at each node from $\mathcal{O}(n_{\mathrm{MAX}}^{I})$ to $\mathcal{O}(n_{\mathrm{MAX}}^{2I})$.
VII-C Multi-Access Design
@@
In this paper we have been focusing on a single multi-hop data link. In practical deployment, a multi-access solution will be mandatory to allow efficient networking, especially in the fronthaul and midhaul domains, where the network exhibits a star topology that each DU is associated with multiple RUs and each CU with multiple DUs. While OFDMA is widely used in modern radio systems and has been dominating most wireless standards, it may not be the optimal match for our DMH-HARQ protocol, due to the challenge in interference management caused by the dynamic traffic pattern that are hard to predict. Similar issue has already been discussed for the original CLARQ problem [33], revealing that time-division multiple access (TDMA) will be a better solution than OFDMA in this context. This will limit the application of the dynamic HARQ approaches such like CLARQ or DMH-HARQ for the air interface between users and base stations, which are generally standardized in modern cellular systems to use OFDMA for better scalability and flexibility. It is also for this reason that we exclude the air interface from the scope of this study.
@@
However, in wireless transport networks which is under our focus in this study, TDMA can still be a good and feasible option, because: i) unlike the air interface of Radio Access Network (RAN), the physical layer design of transport network is not universally standardized but upon the design of vendor, allowing to use TDMA; ii) compared to the number of UEs in RAN, the number of transport network nodes are limited, leading to low demand for scalability; iii) the transport network nodes are generally installed at fixed position, leading to easy synchronization and low demand for flexibility.
@@ Appendix A Proof of Lemma 2
$\begin{split}\xi=&1-\epsilon_{I}(m_{1})+\epsilon_{I}(m_{1})\left[1-\epsilon_{I}(m_{2})\right]\\ &+\epsilon_{I}(m_{1})\epsilon_{I}(m_{2})\left[1-\epsilon_{I}(m_{3})\right]+\dots\\ &+\epsilon_{I}(m_{1})\dots\epsilon_{I}(m_{K-1})\left[1-\epsilon_{I}(m_{K})\right]\\ =&1-\prod\limits_{k=1}^{K}\epsilon_{I}(m_{k})\end{split}$ (15)
@@
$\xi_{\mathrm{os}}=1-\epsilon_{I}\left(\sum\limits_{k=1}^{K}m_{k}\right).$ (16)
@@
$\sum\limits_{k=1}^{K}m_{k}\leqslant\frac{t_{s}-KT_{\mathrm{l}}}{T_{\mathrm{b}}}\leqslant\frac{t_{s}-T_{\mathrm{l}}}{T_{\mathrm{b}}},$ (17)
@@
$\prod\limits_{k=1}^{K}\epsilon_{I}(m)\geqslant\epsilon_{I}\left(\frac{t_{s}-T_{\mathrm{l}}}{T_{\mathrm{b}}}\right),$ (18)
@@ Appendix B Proof of Theorem 1
$n_{\mathrm{sub}}(s)=\begin{cases}n_{\mathrm{opt}}(s^{<-\Delta t>})+\frac{\Delta t}{T_{\mathrm{b}}}&i_{s}=1,\\ n_{\mathrm{opt}}(s^{<-\Delta t>})&\mathrm{otherwise}.\end{cases}$ (19)
@@
$s^{\prime}=\begin{cases}s&i_{s^{\prime}}=0,\\ s^{<\Delta t>}&\text{otherwise}.\end{cases}$ (20)
@@
$\varepsilon^{\prime}_{s^{\prime}}\begin{cases}<\varepsilon_{s}&i_{s^{\prime}}=1,\\ =\varepsilon_{s}&\text{otherwise}.\end{cases}$ (21)
Recalling Eq. (12), it always hold $\xi(\mathfrak{T}^{\mathrm{sub}}_{s_{1}})>\xi_{\mathrm{max}}(s_{0})$. Meanwhile, there is always some $n^{\prime}_{\mathrm{opt}}$ that generates an optimal schedule $\mathfrak{T}^{\mathrm{opt}}_{s_{1}}$ with maximal $\xi_{\mathrm{max}}(s_{1})\geqslant\xi(\mathfrak{T}^{\mathrm{sub}}_{s_{1}})$. Thus:
$\xi_{\mathrm{max}}(s_{1})=\xi_{\mathrm{max}}(s_{0}^{<\Delta t>})>\xi_{\mathrm{max}}(s_{0})\quad\forall s_{0},\forall\Delta t>0,$ (22)
@@
Next, we come to prove the continuity of this function. According to Eq. (8), $\varepsilon_{s}$ is continuous function of $t_{s}$, so $\xi_{\mathrm{max}}(s)$ can be guaranteed $t_{s}$-continuous, if both $\xi_{\mathrm{max}}(s_{\mathrm{l}}(\mathfrak{T}_{s}))$ and $\xi_{\mathrm{max}}(s_{\mathrm{r}}(\mathfrak{T}_{s}))$ are continuous functions of $t_{s}$. This sufficient condition can be derived via mathematical induction as follows.

### v4

@@ VI-C Results and Analyses
Figure 4: Sensitivity tests regarding (4(a)) the time frame length, (4(b)) the number of hops, (4(c)) the decoding/feedback delay, and (4(d)) the pathloss, respectively.
We observe from the results that the DMH-HARQ approach consistently outperforms all baselines, even when under an unfair disadvantage in soft combining. It is promised to hold even more performance gain over the static HARQ solutions when applied with a more advanced combining technology. This excellent performance is credited to the DP algorithm that guarantees to achieve the theoretical maximum of the Bellman sum (6n). Especially, with more available blocklength per hop, since the overall possible numbers of HARQ attempts over the $I$-hop chain increases, this performance gain brought by DP is also increasing. To the contrary, when the channel is harsh or the radio resource is extremely limited, the constraint (6a) is likely forcing to allow only one transmission slot per hop, so that the optimal DMH-HARQ schedule converges to the optimal one-shot-per-hop schedule, which just gives the performance upper bound of optimal static HARQ-IR in FBL [5]. Moreover, the performance gain of DMH-HARQ over the baselines remains significant when the number of hops increases.
Remark that the listening-based cooperative ARQ approach also allocates the radio resource to different hops in a technically dynamic manner. However, its mechanism design only allows the successor hops to inherit unused resource from predecessors, not the other way around. With a same fixed blocklength for every (re)transmission, it results in a longer effective sub-frame for the late hops that are closer to the final information sink than the for early hops that are closer to the source. However, as it can be derived from [39], as a stochastic DP, the optimal solution shall be just the opposite: to allocate more resource to the early hops than the late ones. In the end, the listening-based cooperative ARQ approach cannot hit the optimum of the DP problem (6), and has only a limited gain from the dynamic scheduling. This impact is even more critical in the FBL regime where ARQ/HARQ generally work poor, leading to a significant gap to the performance in the infinite blocklength (IBL) regime. As we see from the results, the listenting-based cooperative ARQ approach performs the worst among all baselines.
@@ VII Further Discussions
VII-A LUT Implementation: Complexity and Performance
As we have discussed in Sec. V-C, the requirement of real-time online solution suggests a LUT-based implementation of the DMH-HARQ method. As the price for the significant time complexity reduction from $\mathcal{O}(n_{\mathrm{MAX}}^{I})$ to $\mathcal{O}\left[\log((n_{\mathrm{MAX}}))\right]$, extra space complexity must be taken into account. As we have already shown, for any given SNR $\gamma$, the allocation policy $n_{\mathrm{opt}}$ is a function of remaining time $t$ and the number of remaining hops $i$. Thus, the size of the LUT implemented at each relaying node is (loosely) upper-bounded by $I_{\mathrm{max}}n_{\mathrm{MAX}}L$, where $n_{\mathrm{MAX}}$ is the maximum number of blocklength in a radio frame, and $L$ is the number of discrete SNR levels. Here we carry out a worst-case analysis for a rough estimate of the memory cost to assess the deployment feasibility.
$I_{\mathrm{max}}$, as aforementioned, is constrained by the E2E latency limit $T_{\mathrm{max}}$, which is typically in the range of tens to hundered milliseconds in our investigated use scenarios. Moreover, for every specific deployment, $I$ is determined by the overall distance of wireless backhaul and the radio coverage of each relay node, which dependents on the frequency band. For long-distance backhaul where rural environments are commonly considered, the traditional sub-$40\text{\,}\mathrm{GHz}$ band is preferred and its typical link length (per hop) is around $15\text{\,}\mathrm{km}$. For midhaul connections in suburban areas, this length reduces to $8\text{\,}\mathrm{km}$ per hop [40]. According to [38], relay chains in midhauls can reach a distance of $40\text{\,}\mathrm{km}$ and in backhauls up to $300\text{\,}\mathrm{km}$, we can reasonably consider that $I_{\mathrm{max}}\leqslant 25$ even in the utmost case. $n_{\mathrm{MAX}}$, on the other hand, depends on the specific radio frame design for the wireless transport network, and is also constrained by the latency limit $T_{\mathrm{max}}$. At last, the SNR quantization levels LLcan be flexibly configured. However, a small $L$ implies higher CSI quantization error, which will certainly degrade the performance. To assess this effect and select the proper $L$, we repeated the benchmark test with default specifications listed in Tab. I, but with $L$-level quantized CSI instead of accurate CSI at each node. The results are illustrated in Fig. 5, which suggests to pick $L\geqslant 64$, so that significant performance loss can be avoided to sufficiently leverage the performance gain of DMH-HARQ.
Given these parameter values, to provide a realistic estimate of the time and memory costs of the LUT implementation, we benchmarked the computation of $L=64$ LUT s regarding Algorithm 1 under different scenarios. The system specifications and the benchmark results are listed in Tab. II, which shows that even for an extremely challenging scenario with $I_{\mathrm{max}}=25$, $n_{\mathrm{MAX}}=5000$, and $d=128$, the offline LUT computation can be accomplished within $5.5$ hours on a regular commercial workstation, and the total size of the generated LUTs is barely above $250\text{\,}\mathrm{kB}$, which is negligible for modern communication devices.
Figure 5: Sensitivity test regarding the CSI quantization levels.
TABLE II: Cost of the offline LUT generation ($L=64$)
CPU Memory
AMD Ryzen 5 5600G @ $3.9\text{\,}\mathrm{GHz}$ $32\text{\,}\mathrm{GB}$
6 cores, 12 threads DDR4-$2133\text{\,}\mathrm{MHz}$
$64\text{\,}\mathrm{bit}$, $7\text{\,}\mathrm{nm}$ process Dual-channel
Operation System LUT Solver
Testing Platform Windows 11 Pro Julia 1.11.6
Scenario Time Memory
$n_{\mathrm{MAX}}=100,~I_{\mathrm{max}}=2,~d=16$ $464.8\text{\,}\mathrm{ms}$ $4.876\text{\,}\mathrm{kB}$
$n_{\mathrm{MAX}}=500,~I_{\mathrm{max}}=4,~d=32$ $31.21\text{\,}\mathrm{s}$ $27.62\text{\,}\mathrm{kB}$
$n_{\mathrm{MAX}}=1000,~I_{\mathrm{max}}=10,~d=64$ $4.021\text{\,}\mathrm{min}$ $48.20\text{\,}\mathrm{kB}$
$n_{\mathrm{MAX}}=5000,~I_{\mathrm{max}}=25,~d=128$ $5.467\text{\,}\mathrm{h}$ $250.88\text{\,}\mathrm{kB}$
VII-B Imperfect CSI Measurement at Local Hop
While we have assumed that each node has perfect CSI at its local hop and statistical CSI at all hops thereafter, it shall be noted that in practical systems, even the CSI measurement at the local hop can be inaccurate or outdated, and some performance degradation will occur in this case, in addition to the impact of the CSI quantization error. Nevertheless, for wireless backhaul and midhaul scenarios where both the transmitter and receiver are mounted at high altitudes with stable positions, the channel variation is generally slow, and the CSI can be accurately estimated with low overhead. Moreover, it is not only our proposed DMH-HARQ scheme that will be affected by such imperfect CSI. Conventional solutions, such like the static HARQ-IR, also rely on CSI to perform optimal blocklength allocation. It is therefore of generic interest to minimize the CSI estimation error and its impact on system performance, which is beyond the scope of this article.
VII-C Timer Overhead
While the DMH-HARQ approach requires a timer field embedded in the message header, it is also worth a discussion to assess its overhead. The timer field must contain sufficient bits to represent the remaining symbols in the current multi-hop forwarding session, so it is lower-bounded by $\lceil\log_{2}(n_{\mathrm{MAX}})\rceil$. We consider the same four scenarios as listed in Tab. II, and calculate the timer-to-payload bit length ratio for each, as shown in Tab. III. Generally, for small data packets over short relay chains, the overhead is significant but acceptable compared to the typical channel coding redundancy, considering the reliability gain it offers. Moreover, this overhead ratio further decreases in long-distance scenarios with larger packets, since the timer grows only logarithmically with $n_{\mathrm{MAX}}$, while decreasing propotionally to the payload length.
TABLE III: Timer-to-payload length ratio of selected scenarios
Scenario Min. Timer Bits Ratio
$n_{\mathrm{MAX}}=100,~d=16$ 4 25.00%
$n_{\mathrm{MAX}}=500,~d=32$ 9 28.13%
$n_{\mathrm{MAX}}=1000,~d=64$ 10 15.63%
$n_{\mathrm{MAX}}=5000,~d=128$ 13 10.16%
VII-D Incremental Redundancy Gain
@@
We have proven that even when applied without any combining technique, DMH-HARQ is able to outperform the optimal static HARQ-IR. A natural idea, of course, is to further enhance the performance of DMH-HARQ with soft combining, especially with an ideal IR, where the PER formula (6i) must be replaced by
$\varepsilon_{s}=Q\left(\sqrt{\frac{n_{s}+\frac{\tau_{s}}{T_{\mathrm{b}}}}{V_{I+1-i_{s}}}}\left(\mathcal{C}_{I+1-i_{s}}-\frac{d}{n_{s}+\frac{\tau_{s}}{T_{\mathrm{b}}}}\right)\ln 2\right).$ (6ax)
@@
Though IR is guaranteed to reduce the achievable E2E packet loss rate of DMH-HARQ, it costs more computational effort. The inclusion of $\tau_{s}$ into $\varepsilon_{s}$ rejects Lemma 1. To keep it an MDP so that the integer DP based framework of Algorithm 1 still applies, $\tau_{s}$ must play its role as part of the system state when solving the decision policy, i.e. $n_{s}$. The dictionaries $\mathcal{D}_{\mathrm{local}}$ and $\mathcal{D}_{\mathrm{succ}}$ in Algorithm 1 shall map $t_{s},i_{s},\tau_{s}]$ instead of $t_{s},i_{s}]$ to $\xi_{\mathrm{opt}},n_{\mathrm{opt}}]$. This will significantly increase the effective size of the state space, raising the time complexity of solving $n_{s}$ at each node from $\mathcal{O}(n_{\mathrm{MAX}}^{I})$ to $\mathcal{O}(n_{\mathrm{MAX}}^{2I})$.
VII-E Multi-Access Design
@@
In this paper we have been focusing on a single multi-hop data link. In practical deployment, a multi-access solution will be mandatory to allow efficient networking, especially in the fronthaul and midhaul domains, where the network exhibits a star topology that each O-DU is associated with multiple O-RUs and each O-CU with multiple O-DUs. While OFDMA is widely used in modern radio systems and has been dominating most wireless standards, it may not be the optimal match for our DMH-HARQ protocol, due to the challenge in interference management caused by the dynamic traffic pattern that are hard to predict. Similar issue has already been discussed for the original CLARQ problem [36], revealing that time-division multiple access (TDMA) will be a better solution than OFDMA in this context. This will limit the application of the dynamic HARQ approaches such like CLARQ or DMH-HARQ for the air interface between users and base stations, which are generally standardized in modern cellular systems to use OFDMA for better scalability and flexibility. It is also for this reason that we exclude the air interface from the scope of this study.
@@
However, in wireless transport networks which is under our focus in this study, TDMA can still be a good and feasible option, because: i) unlike the air interface of Radio Access Network (RAN), the physical layer design of transport network is not universally standardized but upon the design of vendor, allowing to use TDMA; ii) compared to the number of user equipments (UEs) in RAN, the number of transport network nodes are limited, leading to low demand for scalability; iii) the transport network nodes are generally installed at fixed position, leading to easy synchronization and low demand for flexibility.
@@ Appendix A Proof of Lemma 2
$\begin{split}\xi=&1-\epsilon_{I}(m_{1})+\epsilon_{I}(m_{1})\left[1-\epsilon_{I}(m_{2})\right]\\ &+\epsilon_{I}(m_{1})\epsilon_{I}(m_{2})\left[1-\epsilon_{I}(m_{3})\right]+\dots\\ &+\epsilon_{I}(m_{1})\dots\epsilon_{I}(m_{K-1})\left[1-\epsilon_{I}(m_{K})\right]\\ =&1-\prod\limits_{k=1}^{K}\epsilon_{I}(m_{k})\end{split}$ (6ay)
@@
$\xi_{\mathrm{os}}=1-\epsilon_{I}\left(\sum\limits_{k=1}^{K}m_{k}\right).$ (6az)
@@
$\sum\limits_{k=1}^{K}m_{k}\leqslant\frac{t_{s}-KT_{\mathrm{l}}}{T_{\mathrm{b}}}\leqslant\frac{t_{s}-T_{\mathrm{l}}}{T_{\mathrm{b}}},$ (6ba)
@@
$\prod\limits_{k=1}^{K}\epsilon_{I}(m)\geqslant\epsilon_{I}\left(\frac{t_{s}-T_{\mathrm{l}}}{T_{\mathrm{b}}}\right),$ (6bb)
@@ Appendix B Proof of Theorem 1
$n_{\mathrm{sub}}(s)=\begin{cases}n_{\mathrm{opt}}(s^{<-\Delta t>})+\frac{\Delta t}{T_{\mathrm{b}}}&i_{s}=1,\\ n_{\mathrm{opt}}(s^{<-\Delta t>})&\mathrm{otherwise}.\end{cases}$ (6bc)
@@
$s^{\prime}=\begin{cases}s&i_{s^{\prime}}=0,\\ s^{<\Delta t>}&\text{otherwise}.\end{cases}$ (6bd)
@@
$\varepsilon^{\prime}_{s^{\prime}}\begin{cases}<\varepsilon_{s}&i_{s^{\prime}}=1,\\ =\varepsilon_{s}&\text{otherwise}.\end{cases}$ (6be)
Recalling Eq. (6n), it always hold $\xi(\mathfrak{T}^{\mathrm{sub}}_{s_{1}})>\xi_{\mathrm{max}}(s_{0})$. Meanwhile, there is always some $n^{\prime}_{\mathrm{opt}}$ that generates an optimal schedule $\mathfrak{T}^{\mathrm{opt}}_{s_{1}}$ with maximal $\xi_{\mathrm{max}}(s_{1})\geqslant\xi(\mathfrak{T}^{\mathrm{sub}}_{s_{1}})$. Thus:
$\xi_{\mathrm{max}}(s_{1})=\xi_{\mathrm{max}}(s_{0}^{<\Delta t>})>\xi_{\mathrm{max}}(s_{0})\quad\forall s_{0},\forall\Delta t>0,$ (6bf)
@@
Next, we come to prove the continuity of this function. According to Eq. (6i), $\varepsilon_{s}$ is continuous function of $t_{s}$, so $\xi_{\mathrm{max}}(s)$ can be guaranteed $t_{s}$-continuous, if both $\xi_{\mathrm{max}}(s_{\mathrm{l}}(\mathfrak{T}_{s}))$ and $\xi_{\mathrm{max}}(s_{\mathrm{r}}(\mathfrak{T}_{s}))$ are continuous functions of $t_{s}$. This sufficient condition can be derived via mathematical induction as follows.